I am trying to relearn some algebra for college and I cant figure out some of the questions... please help
-3/5[2/3-1/2(3/4- 1/2)]
can you break it down for me?

Answer by rfer(16322)

-3/5[2/3-1/2(3/4-1/2)]
just follow the oeder of operations.
-3/5[2/3-1/2(1/4)]
-3/5[2/3-1/8]
-3/5[16/24-3/24]
-3/5*13/24=-39/120=-13/40